OpenStudy (xenesis98):

If A and B are dependent events, which of these conditions must be true? a. P(A and B) = P(A) + P(B) b. P(A and B) = P(B) P(A) c. P(B|A) = P(B) d. P(A|B) = P(A) e. P(B|A) ≠ P(B)

satellite73 (satellite73):

a bunch of them

satellite73 (satellite73):

the definition is this A and B are independent if \[P(A|B)=P(A)\] the left hand side of the equal sign is \[\frac{P(A\cap B)}{P(B)}\]so it also says \[\frac{P(A\cap B)}{P(B)}=P(A)\]

satellite73 (satellite73):

you can change this equality around using algebra

OpenStudy (xenesis98):

so D.?

OpenStudy (reemii):

not only. satellite73 gave a hint.

OpenStudy (xenesis98):

oh ok